Please see below help would be appreciated, thanks in advance!
Answer #1

a) v^2 = v0^2 + 2 a x

5^2 = 0^2 + 2*a*2

25 = 4 a

a = 6.25 m/s^2

b) F = ma = 2*6.25 = 12.5 N

c)

now F - friction = m a

friction = u m g

F = um g + ma = 0.5*2*9.81+12.5=22.31 N

d)same as part a

a = 6.25 m/s^2

e) now F - mg sin theta = m a

F = mg sin theta + m a = 2*9.81*sin(30 degrees) +12.5= 22.31 N

f) so sum forces perp to slope

N - m g cos theta = 0

N = mg cos theta
so friction = u N = u m g cos theta

F - m g sin theta - friction = ma

F = 2*9.81*sin(30 degrees) +12.5 + 0.5*2*9.81*cos(30 degrees)= 30.8 n

g)
now when you sum perp to slope

N - mg cos theta - F sin theta = 0

N = m g cos theta + F sin theta

so now sum with slope

F cos theta - m g sin theta - u ( mg cos theta + F sin theta) = ma


F*cos(30 degrees) - 2*9.81*sin(30 degrees) - 0.5*(2*9.81*cos(30 degrees) +F*sin(30 degrees) ) = 2*6.25

F=50 N
Answer #2

a)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

b)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/

Fnet=F=ma=2*6.25=12.5N

c)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

Fnet=ma=2*6.25=12.5N

F-f=Fnet

F=Fnet+f=12.5+umg=12.5+0.5*2*9.81=22.31N

d)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

e)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

Fnet=ma=2*6.25=12.5N

F-mgsin(theta)=Fnet

F=Fnet+mgsin(theta)=12.5+2*9.81*sin(30)=12.5+2*9.81*0.5=22.31N

f)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

F-mgsin(theta)-f=Fnet

F=Fnet+mgsin(theta)+f=Fnet+mgsin(theta)+umgcos(theta)=12.5+2*9.81*sin(30)+0.5*2*9.81*cos(30)=

=12.5+2*9.81*0.5+0.5*2*9.81*0.866=30.8057N

g)

V^2=U^2+2aS

25=0+2a*2

a=6.25m/s2

Fnet=ma=2*6.25=12.5N

Fcos30-mgsin(theta)-f=Fnet

Fcos(30)=Fnet+mgsin(theta)+f=Fnet+mgsin(theta)+u(mgcos(theta)+Fsin(30))

0.866F=12.5+2*9.81*sin(30)+0.5(2*9.81*0.866+0.5F)

0.366F=12.5+2*9.81*0.5+0.5*2*9.81*0.866=30.8057N

F=30.8057/0.366=84.1686N